OpenStudy (anonymous):

a woman stands on the bank of a frozen lake with a dog by her side. she skims a bone across the ice at a speed of 3ms. the bone slows down with deceleration 0.4 ms square and the dog chases it with acce;eration 0.6 ms square. how far out from the bank does the dog catch up with the bone

OpenStudy (anonymous):

the equation describing the motion of the bone is : xbone = 3t - 0.2t^2 the equation describing the motion of the dog it xdog = 0.3t^2 xbone=xdog (when the meet) 3t - 0.2t^2 = 0.3t^2 3t -0.5t^2 = 0 0.5t ( 6 - t) = 0 t = 6 plugging 6 in either xbone or xdog gives you : 10.8